OpenStudy (anonymous):

A firecracker shoots up from a hill 160 feet high, with an initial speed of 90 feet per second. Using the formula H(t) = -16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? Medal and Fan!! Help please...

OpenStudy (campbell_st):

well after substituting you get \[H(t) = -16t^2 + 90t + 160\] to find the time when it hits the ground set H(t) = 0 and then solve for t. Only consider the positive time value. I think you need to general quadratic formula.

OpenStudy (anonymous):

I got 8 but some how i feel thats wrong can u walk me through the equation?

OpenStudy (campbell_st):

well you can divide very term by 2 \[0 = -8t^2 + 45t + 80\] then \[t = \frac{-45 \pm \sqrt{45^2 - 4 \times -8 \times 80}}{2 \times -8}\] \[t = \frac{-45 \pm \sqrt{4585}}{-32}\] hope that helps

OpenStudy (anonymous):

so it is 8...

OpenStudy (campbell_st):

oops should be \[t = \frac{45 + \sqrt{4585}}{16}\] the denominator should be -16 not -32 and I didn't get 8

OpenStudy (anonymous):

ok I just re did it and I got 6 ... please tell me im right...@campbell_st

OpenStudy (campbell_st):

well \[\sqrt{4585} = 67.7\] so solve \[t = \frac{45 + 67.7}{16}\]

OpenStudy (anonymous):

thanks for the help I got it ^.^